You have $10,000 to invest in a stock portfolio. Your choices are Stock X with a...

  1. You have $10,000 to invest in a stock portfolio. Your choices are Stock X with a return of 15 percent and Stock Y with a return of 10 percent. If your goal is to create a portfolio with a return of 13.5 percent, how much money will you invest in Stock X? In Stock Y?

PLEASE GIVE ME A FULL EXPLANATION WITH ALL THE ANSWERS, THANK YOU
Answer #1

Let investment in X=$x

Hence investment in Y=(10,000-x)

Portfolio return=Respective return*Respective weight

13.5=(x/10,000*15)+(10,000-x)/10,000*10

(13.5*10,000)=15x+100,000-10x

135,000=15x+100,000-10x

x=(135,000-100,000)/(15-10)

=$7000=investment in X

Hence investment in Y=(10,000-x)

=$3000